# CREW: Facilitating Human-AI Teaming Research

### v1

@@ Abstract
With the increasing deployment of artificial intelligence (AI) technologies, the potential of humans working with AI agents has been growing at a great speed. Human-AI teaming is an important paradigm for studying various aspects when humans and AI agents work together. The unique aspect of Human-AI teaming research is the need to jointly study humans and AI agents, demanding multidisciplinary research efforts from machine learning to human-computer interaction, robotics, cognitive science, neuroscience, psychology, social science, and complex systems. However, existing platforms for Human-AI teaming research are limited, often supporting oversimplified scenarios and a single task, or specifically focusing on either human-teaming research or multi-agent AI algorithms. We introduce CREW, a platform to facilitate Human-AI teaming research and engage collaborations from multiple scientific disciplines, with a strong emphasis on human involvement. It includes pre-built tasks for cognitive studies and Human-AI teaming with expandable potentials from our modular design. Following conventional cognitive neuroscience research, CREW also supports multimodal human physiological signal recording for behavior analysis. Moreover, CREW benchmarks real-time human-guided reinforcement learning agents using state-of-the-art algorithms and well-tuned baselines. With CREW, we were able to conduct 50 human subject studies within a week to verify the effectiveness of our benchmark.
@@ 1 Introduction
Unlike traditional AI research, which typically focuses on machine learning algorithms in isolation, Human-AI teaming requires a multidisciplinary approach to incorporate insights from various scientific domains. Numerous studies have examined human-human teaming [29] with cognitive science, neuroscience, and psychology. Machine learning and robotics communities have extensively researched multi-agent machine learning [61], while team dynamics [44] has been explored in complex systems, social science, and network science. Despite the importance and potential of this research paradigm, there is still a lack of a comprehensive and unified platform to benefit research on joint efforts across disciplines and scalable hypothesis verification.
Developing a comprehensive platform for Human-AI teaming research presents several unique challenges. Firstly, the platform needs to support diverse tasks with varying complexities with easy extensions for new tasks or modifications. While reinforcement learning research platforms [51] have widely adopted this practice, current Human-AI teaming platforms remain limited to single tasks [13, 54]. Secondly, enabling real-time communication through various modalities between multiple humans and heterogeneous AI policies is essential for effective collaboration. However, existing solutions typically support human feedback only through replaying offline trajectories and do not implement real-time feedback mechanisms. Understanding how to build AI that can team with, learn from, be guided by, align with, and augment humans is as crucial as modeling human behavior during interactions with AI. Therefore, the platform must provide interfaces for recording human physiological data alongside agent data, tailored for cognitive science and neuroscience studies. Furthermore, current studies often involve small participant groups, making it difficult to derive generalizable conclusions. Lastly, the absence of a unified platform has limited fully open-sourced studies and the establishment of appropriate benchmarks.
@@ 2 Related Work
Human-AI Teaming Research Extensive research has explored Human-AI teaming across various domains. Machine learning studies have developed algorithms to leverage human expertise to improve the accuracy [33], robustness [8, 63], and interpretability [33, 8, 62] of models. Integrating human feedback can not only improve performance [27, 18] but also align the models with human preference [33, 40, 63]. Human-computer interaction research has created interfaces [56, 37] and workflows [9] that enhance collaboration between humans and AI, combining their strengths to achieve superior performance. Ethical research focuses on understanding and mitigating the societal [24, 17], ethical [19, 59, 43], and technical [22, 49] challenges of the rapid advancement and wide adoption of AI. Many fields, including neuroscience, healthcare, robotics, transportation, education, security, and accessibility, have shown growing interest [38, 25, 6, 39, 42, 32] in Human-AI teaming. Overall, the broad spectrum of interests highlights the need for multidisciplinary collaboration to drive further advancements.
Human-AI Teaming Platform While significant progress has been made in Human-AI teaming research, there remains an absence of a comprehensive research platform. Overcooked Environment [13] is a simplified version of the original popular game to challenge human agents and AI agents in tasks that require close coordination and strategic teamwork. StarCraft II Learning Environment (SC2LE) [54] supports adversarial settings to allow Human-AI interaction and learning from human demonstrations. Rapid Integration and Development Environment(RIDE) [52] focuses on defense-related scenarios, emphasizing rapid development and integration of AI systems for operational purposes. In addition to real-time decision-making tasks, previous research has developed platforms [21, 36, 60] that focus on offline preference or rating settings where humans can provide offline evaluations or corrections with imitation learning or reinforcement learning.
Table 1: Human-AI teaming platforms
@@
Overcooked [13] 1+ 1
SC2LE [54] 1 v 1
RIDE [52] No Limit
@@
Table 2: *Instead of real-time feedback training, RLHF-Blender and Uni-RLHF support online training where humans are presented with data from the replay buffer instead of the current experience.
However, existing platforms have more than one of the following limitations that constrain Human-AI teaming research as summarized in Tab. 2. Most environments only support one type of task that can be difficult to extend, and the interactions between humans and machine learning agents are limited to mouse and keyboard operations. Moreover, most of the environments only support the data collection on the game data, such as state, action, or reward focusing on the machine learning algorithm development, but none of them support the collection and analysis of human physiological data (eye movement, pupillometry, brain activity, heart impulse, or natural language) to understand human cognitive states and different effects on human subjects. Furthermore, the scale of collaboration has been limited to two agents in a collaborative or competitive setting. Notably, most existing Human-AI teaming studies have only been evaluated on a very small number of subjects or among the authors, which greatly limits our understanding of the state-of-the-art performance. There remains a large gap between the existing platforms and the desired environment to facilitate future research.
Real-Time Human-Guided Learning Most real-world decision-making processes require rapid decisions and adaptation in real time. Therefore, real-time human-guided learning is an essential topic in Human-AI teaming research. Previous research has focused on algorithm design to integrate real-time human feedback into policy training. TAMER and Deep TAMER [30, 55] learn to predict human discrete feedback and utilize the feedback model as the value function for policy learning. COACH and DeepCOACH [5, 35] addressed the inconsistency of human feedback by modeling it as the advantage function under an actor-critic framework. Recent progress has further refined these algorithms, addressing various challenges in human feedback integration, such as different feedback modalities  feedback stochasticity [4], and continuous action spaces [46]. Other works have explored indirectly inferring human objectives from feedback or preference [26, 57, 18]. CREW provides an extensive environment for real-time human-guided learning with online training and feedback integration, human physiological data collection, and parallel distributed experiment support.
@@ 3.1 Platform Vision and Design Philosophy
We design CREW to facilitate Human-AI teaming research. Our vision is to construct a unified platform for researchers from diverse backgrounds, allowing them to join forces from human-AI interaction, machine learning algorithms, workflow design, as well as human analysis and training. To achieve this, CREW incorporates the following capabilities, as illustrated in Fig. 1.
Extensible and open environment design. CREW provides built-in tasks for rapid development and allows users to integrate customized tasks to accommodate the limitless applications of Human-AI teaming.
@@ 3.2 Environments
Tasks We select Unity as the simulation engine for CREW due to its popularity in game design and AI research to allow extensible and open environment design. We have implemented four challenging tasks as examples. Multi-player tasks are designed for multi-agent and multi-human teaming research, and single-player tasks are designed for human-guided AI agent learning studies. For each task, we provide both visual and structured state input options. The detailed settings are summarized in Tab. 3.
@@
is a modified version of Atari bowling where each round consists of 10 rolls and the score for each roll is the number of pins hit. Bowling is designed to have the simplest dynamics among our tasks to serve as a rapid test for training a single agent. Find Treasure (Fig. 2A) is a single-player embodied navigation task where the agent’s goal is to explore a maze and reach the treasure with randomized initial and goal locations. 1v1 Hide-and-Seek [14, 15] (Fig. 2B) is a one-on-one hide-and-seek task where the seeker learns to explore the maze and catch a moving hider that follows a heuristic policy for obstacle avoidance, and run away from the seeker within line of sight. We introduce this task as an adversarial competition setting. NvN Hide-and-Seek (Fig. 2C) is a multiplayer version where multiple seekers and hiders can coordinate, collaborate, and compete. The hiders and seekers can either be controlled by humans or heterogeneous AI policies.
@@
Visual Observations Different visual observations create various challenges in visual embodiment learning for both humans [53, 20] and AI agents [7, 50]. We provide various camera configurations to support perceptual-motor research. As shown in Fig 2, for all our navigation and competitive tasks, we implemented visual observations from multiple views for the users’ selection: a top-down fully observable view, a top-down accumulated partially observable view similar to SLAM in robotics [48], a top-down egocentric view, and a first-person view.
@@
Procedural Generation Learning robust, generalizable, and scalable AI agents requires diverse training environments. Procedural generation allows for the creation of a wide range of environment patterns and terrain types. We provide randomized maze patterns where the grids are guaranteed to be connected (Fig. 3A) and procedural-generated terrains for more complex simulations as in Fig. 3B.
@@ 3.3 Human and Agent Role Assignment
Human Feedback Interface We provide a user interface to allow the Viewer role to provide direct feedback to AI agents shown in Fig. 4. Scalar feedback is the most common feedback type used in human-guided RL [30, 55, 57]. Conventionally, a human chooses to provide a positive or negative rating to the state-action pair of an agent policy. Additionally, CREW offers an interface that allows humans to provide feedback that is continuous in value and time, enabling granular feedback information. Moreover, our interface also allows humans to directly take control over agents and perform teleoperation.
@@ 3.4 Multiplayer and Parallel Sessions
Enabling multi-human multi-agent sessions requires robust networking solutions (Fig. 5). We use Unity Netcode [2] for game state synchronization, and Nakama [1] as the networking server. In CREW, a server instance hosts the task, runs the simulation, and handles agent policy training, which can be executed on a powerful headless GPU server. Human participants can connect via client instances on less powerful machines, which display synchronized game states and collect human input. CREW is cross-platform, allowing participation from Linux, Windows or MacOS machines.
@@ 3.5 Human and Agent Data Collection
Human data Besides the feedback interfaces that collect feedback signals of multiple modalities and teleoperation actions, we also provide interfaces for collecting audio, eye gaze, pupillometry, electroencephalogram (EEG), and electrocardiogram (ECG) physiological responses as in Fig. 6.
Agent data including the policy weights, observations, actions, rewards, feedback received, and loss values at every time step can all be saved for further analysis. Users also have the option to enable experiment monitoring and logging by Weights & Biases [3]. As all of our tasks include vision-based settings, we also provide implementations of a set of vision encoder architectures.
Figure 6: Data collection in CREW. All human and agent data, including game states, feedback, mouse and keyboard, audio, gaze and pupil data, EEG and ECG data are all streamed and synchronized through Lab Streaming Layer [31].
@@ 3.6 Designing Algorithms
Algorithms research is crucial for Human-AI teaming. We designed the algorithm component of CREW to be extensible and accessible to the ML community. Algorithms are implemented in Python with PyTorch [41] and TorchRL [12] which is part of the PyTorch ecosystem, making it easy for researchers to design and deploy new algorithms. We provide implementations of state-of-the-art human-guided machine learning algorithms and strong reinforcement learning baselines. The API for communication between a Unity instance and Python algorithm is implemented with MLAgents [28]. All environments follow a uniform communication protocol for the observation and action data across tasks. Our modular design allows smooth switching between tasks and algorithms. Real human experiments are time-consuming and costly. To ease algorithm debugging, we implemented simulated human feedback providers as surrogates for real humans for complex tasks. These simulated feedback uses heuristics based on prior task knowledge and is not available in novel tasks in the real world; hence, they should be used solely as debugging tools, not as replacements for real human evaluations.

### v3

@@ Abstract
With the increasing deployment of artificial intelligence (AI) technologies, the potential of humans working with AI agents has been growing at a great speed. Human-AI teaming is an important paradigm for studying various aspects when humans and AI agents work together. The unique aspect of Human-AI teaming research is the need to jointly study humans and AI agents, demanding multidisciplinary research efforts from machine learning to human-computer interaction, robotics, cognitive science, neuroscience, psychology, social science, and complex systems. However, existing platforms for Human-AI teaming research are limited, often supporting oversimplified scenarios and a single task, or specifically focusing on either human-teaming research or multi-agent AI algorithms. We introduce CREW, a platform to facilitate Human-AI teaming research in real-time decision-making scenarios and engage collaborations from multiple scientific disciplines, with a strong emphasis on human involvement. It includes pre-built tasks for cognitive studies and Human-AI teaming with expandable potentials from our modular design. Following conventional cognitive neuroscience research, CREW also supports multimodal human physiological signal recording for behavior analysis. Moreover, CREW benchmarks real-time human-guided reinforcement learning agents using state-of-the-art algorithms and well-tuned baselines. With CREW, we were able to conduct 50 human subject studies within a week to verify the effectiveness of our benchmark.
@@ 1 Introduction
Unlike traditional AI research, which typically focuses on machine learning algorithms in isolation, Human-AI teaming requires a multidisciplinary approach to incorporate insights from various scientific domains. Numerous studies have examined human-human teaming Klimoski & Mohammed (1994) with cognitive science, neuroscience, and psychology. Machine learning and robotics communities have extensively researched multi-agent machine learning Zhang et al. (2021), while team dynamics Salas et al. (2018) has been explored in complex systems, social science, and network science. Despite the importance and potential of this research paradigm, there is still a lack of a comprehensive and unified platform to benefit research on joint efforts across disciplines and scalable hypothesis verification.
Developing a comprehensive platform for Human-AI teaming research presents several unique challenges. Firstly, the platform needs to support diverse tasks with varying complexities with easy extensions for new tasks or modifications. While reinforcement learning research platforms Towers et al. (2023) have widely adopted this practice, current Human-AI teaming platforms remain limited to single tasks Carroll et al. (2019); Vinyals et al. (2017). Secondly, enabling real-time communication through various modalities between multiple humans and heterogeneous AI policies is essential for effective collaboration. However, existing solutions typically support human feedback only through replaying offline trajectories and do not implement real-time feedback mechanisms. Understanding how to build AI that can team with, learn from, be guided by, align with, and augment humans is as crucial as modeling human behavior during interactions with AI. Therefore, the platform must provide interfaces for recording human physiological data alongside agent data, tailored for cognitive science and neuroscience studies. Furthermore, current studies often involve small participant groups, making it difficult to derive generalizable conclusions. Lastly, the absence of a unified platform has limited fully open-sourced studies and the establishment of appropriate benchmarks.
@@ 2 Related Work
Human-AI Teaming Research Extensive research has explored Human-AI teaming across various domains. Machine learning studies have developed algorithms to leverage human expertise to improve the accuracy Lertvittayakumjorn et al. (2020), robustness Bao et al. (2018); Ziegler et al. (2019), and interpretability Lertvittayakumjorn et al. (2020); Bao et al. (2018); Zhou & Chen (2019) of models. Integrating human feedback can not only improve performance Ibarz et al. (2018); Christiano et al. (2017) but also align the models with human preference Lertvittayakumjorn et al. (2020); Ouyang et al. (2022); Ziegler et al. (2019). Human-computer interaction research has created interfaces Weisz et al. (2021); Neerincx et al. (2018) and workflows Bau et al. (2020) that enhance collaboration between humans and AI, combining their strengths to achieve superior performance. Ethical research focuses on understanding and mitigating the societal Hemmer et al. (2023); Chowdhury et al. (2022), ethical Ezer et al. (2019); Yin et al. (2019); Pflanzer et al. (2023), and technical Green & Chen (2019); Stahl et al. (2021) challenges of the rapid advancement and wide adoption of AI. Many fields, including neuroscience, healthcare, robotics, transportation, education, security, and accessibility, have shown growing interest Nourbakhsh et al. (2005); Henry et al. (2022); Atakishiyev et al. (2021); Nwana (1990); Pazho et al. (2023); Kumar & Jain (2022) in Human-AI teaming. Overall, the broad spectrum of interests highlights the need for multidisciplinary collaboration to drive further advancements.
Human-AI Teaming Platform While significant progress has been made in Human-AI teaming research, there remains an absence of a comprehensive research platform. Overcooked Environment Carroll et al. (2019) is a simplified version of the original popular game to challenge human agents and AI agents in tasks that require close coordination and strategic teamwork. StarCraft II Learning Environment (SC2LE) Vinyals et al. (2017) supports adversarial settings to allow Human-AI interaction and learning from human demonstrations. Rapid Integration and Development Environment(RIDE) USC Institute for Creative Technologies (2024) focuses on defense-related scenarios, emphasizing rapid development and integration of AI systems for operational purposes. In addition to real-time decision-making tasks, previous research has developed platforms Freire et al. (2020); Metz et al. (2023); Yuan et al. (2024) that focus on offline preference or rating settings where humans can provide offline evaluations or corrections with imitation learning or reinforcement learning.
@@
Overcooked Carroll et al. (2019) 1+ 1
SC2LE Vinyals et al. (2017) 1 v 1
RIDE USC Institute for Creative Technologies (2024) No Limit
@@
Table 1: Human-AI teaming platforms. *Instead of real-time feedback training, RLHF-Blender and Uni-RLHF support online training where humans are presented with data from the replay buffer instead of the current experience.
Overall, existing platforms have more than one of the following limitations that constrain Human-AI teaming research as summarized in Table. 1. Most environments only support one type of task that can be difficult to extend, and the interactions between humans and machine learning agents are limited to mouse and keyboard operations. Moreover, most of the environments only support collecting game data, such as state, action, or reward focusing on the machine learning algorithm development, but none of them support the collection and analysis of human physiological data (eye movement, pupillometry, brain activity, heart impulse, or natural language) to understand human cognitive states and different effects on human subjects, as often required in the cognitive studies involved in Human-AI teaming Thakur et al. ; Dikker et al. (2017); Heinisch et al. (2024); Gordon et al. (2020); Koorathota et al. (2023); Xu et al. (2021). Additionally, the scale of collaboration has been limited to two agents in a collaborative or competitive setting. Notably, most existing Human-AI teaming studies have only been evaluated on a very small number of subjects or among the authors ($N<10$, mostly $N<5$), which greatly limits our understanding of the state-of-the-art performance. Furthermore, common previous multi-agent environments such as OvercookedCarroll et al. (2019) and Starcraft Vinyals et al. (2017) are closed-sourced, making it unrealistic to build additional features on top of them. This also makes it impossible for researchers to modify the tasks to their own needs. There remains a large gap between the existing platforms and the desired environment to facilitate future research.
Real-Time Human-Guided Learning Most real-world decision-making processes require rapid decisions and adaptation in real time. Therefore, real-time human-guided learning is an essential topic in Human-AI teaming research. Previous research has focused on algorithm design to integrate real-time human feedback into policy training. TAMER and Deep TAMER Knox & Stone (2009); Warnell et al. (2018) learn to predict human discrete feedback and utilize the feedback model as the one-step myopic value function for policy learning. COACH and DeepCOACH Arumugam et al. (2019); MacGlashan et al. (2016) addressed the inconsistency of human feedback by modeling it as the advantage function under an actor-critic framework. Recent progress has further refined these algorithms, addressing various challenges in human feedback integration, such as different feedback modalities  feedback stochasticity Arakawa et al. (2018), and continuous action spaces Sheidlower et al. (2022). Other works have explored indirectly inferring human objectives from feedback or preference Huang et al. (2023); Xiao et al. (2020); Christiano et al. (2017). CREW provides an extensive environment for real-time human-guided learning with online training and feedback integration, human physiological data collection, and parallel distributed experiment support.
@@ 3.1 Platform Vision and Design Philosophy
We design CREW to facilitate Human-AI teaming research. Our vision is to construct a unified platform for researchers from diverse backgrounds, allowing them to join forces from human-AI interaction, machine learning algorithms, workflow design, as well as human analysis and training. To achieve this, CREW incorporates the following capabilities, as illustrated in Figure. 1.
Extensible and open environment design. CREW is fully open-sourced. Our key contribution is building the infrastructure to allow the development of scalable environments with supported Human-AI research features that scale up to very complex settings. We implemented a set of built-in example tasks described in Section 3.2 for rapid development. Many additional features and game logic can be added without modifying complex infrastructure-level components.
@@ 3.2 Environments
Tasks We select Unity as the simulation engine for CREW due to its popularity in game design and AI research to allow extensible and open environment design. Unity, as used by many popular commercial games as has a high ceiling for the complexity of game development. It also has a large user community which makes customized task development more accessible. Adding new components to an environment can be done through simple drag-and-drop (on the left of Figure. 2 ).
We have implemented four challenging tasks as examples. Multi-player tasks are designed for multi-agent and multi-human teaming research, and single-player tasks are designed for human-guided AI agent learning studies. For each task, we provide both visual and structured state input options. The detailed settings are summarized in Table. 2.
@@
is a modified version of Atari bowling where each round consists of 10 rolls and the score for each roll is the number of pins hit. Bowling is designed to have the simplest dynamics among our tasks to serve as a rapid test for training a single agent. Find Treasure (Figure. 3A) is a single-player embodied navigation task where the agent’s goal is to explore a maze and reach the treasure with randomized initial and goal locations. 1v1 Hide-and-Seek Chen et al. (2020; 2021a) (Figure. 3B) is a one-on-one hide-and-seek task where the seeker learns to explore the maze and catch a moving hider that follows a heuristic policy for obstacle avoidance, and run away from the seeker within line of sight. We introduce this task as an adversarial competition setting. NvN Hide-and-Seek (Figure. 3C) is a multiplayer version where multiple seekers and hiders can coordinate, collaborate, and compete. The hiders and seekers can either be controlled by humans or heterogeneous AI policies.
@@
Customizing new tasks or easily extend current tasks. Researchers can build tasks of much higher flexibility and complexity in CREW than existing platforms due to our efforts in setting distributed multi-agent systems mixing human and AI agents. We chose the set of tasks above only for benchmarking purposes for human-guided reinforcement learning experiments, and it is far below the complexity limit of game design that CREW can support. As shown in Figure. 2 on the right side, we can easily scale up the difficulty of a Hide-and-Seek game to a more complex map. Another example is a recent work Ji et al. (2024) that shows the usage of CREW for human-guided multi-agent Hide-and-Seek.
Figure 2: Left: Example of adding objects to an environment in CREW through drag-and-drop. Right: Example of scaling up the complexity of hide-and-seek to a search-and-rescue task.
@@
Visual Observations Different visual observations create various challenges in visual embodiment learning for both humans Vander Heyden et al. (2017); Fishbein et al. (1972) and AI agents Bandini & Zariffa (2020); Tarun et al. (2019). We provide various camera configurations to support perceptual-motor research. As shown in Figure.3, for all our navigation and competitive tasks, we implemented visual observations from multiple views for the users’ selection: a top-down fully observable view, a top-down accumulated partially observable view similar to SLAM in robotics Smith & Cheeseman (1986), a top-down egocentric view, and a first-person view.
@@
Procedural Generation Learning robust, generalizable, and scalable AI agents requires diverse training environments. Procedural generation allows for the creation of a wide range of environment patterns and terrain types. We provide randomized maze patterns where the grids are guaranteed to be connected (Figure. 4A) and procedural-generated terrains for more complex simulations as in Figure. 4B.
@@ 3.3 Human and Agent Role Assignment
Human Feedback Interface We provide a user interface to allow the Viewer role to provide direct feedback to AI agents shown in Figure. 5. Scalar feedback is the most common feedback type used in human-guided RL Knox & Stone (2009); Warnell et al. (2018); Xiao et al. (2020). Conventionally, a human chooses to provide a positive or negative rating to the state-action pair of an agent policy. Additionally, CREW offers an interface that allows humans to provide feedback that is continuous in value and time, enabling granular feedback information. Moreover, our interface also allows humans to directly take control over agents and perform teleoperation.
@@ 3.4 Multiplayer and Parallel Sessions
Enabling multi-human multi-agent sessions requires robust networking solutions (Figure. 6). We use Unity Netcode net for game state synchronization, and Nakama nak as the networking server. In CREW, a server instance hosts the task, runs the simulation, and handles agent policy training, which can be executed on a powerful headless GPU server. Human participants can connect via client instances on less powerful machines, which display synchronized game states and collect human input. CREW is cross-platform, allowing participation from Linux, Windows or MacOS machines.
@@ 3.5 Human and Agent Data Collection
Human data Prior research in Human-AI teaming and human-human teaming has leveraged physiological features in various ways. Thakur et al. uses predicting human eye gaze as an auxiliary task for guiding imitation learning. Dikker et al. (2017) conducts brain synchrony with EEG in a classroom multi-human teaming setup. Qin et al. (2022) leverages pupil dynamics to predict the team performance in a multi-human virtual reality task. Heinisch et al. (2024) uses physiological data for human-robot interaction with service robots. Gordon et al. (2020) explored physiological synchrony in teaming. Koorathota et al. (2023) used gaze data to enhance vision transformers. Xu et al. (2021) uses EEG data as implicit human feedback for accelerating RL. We implemented the relevant features to support future research along these lines. Besides the feedback interfaces that collect feedback signals of multiple modalities and teleoperation actions, we also provide interfaces for collecting audio, eye gaze, pupillometry, electroencephalogram (EEG), and electrocardiogram (ECG) physiological responses as in Figure. 7 with time synchronizations from each machine with Lab Streaming Layer Kothe et al. (2024).
Agent data including the policy weights, observations, actions, rewards, feedback received, and loss values at every time step can all be saved for further analysis. Users also have the option to enable experiment monitoring and logging by Weights & Biases Biewald (2020). As all of our tasks include vision-based settings, we also provide implementations of a set of vision encoder architectures.
Figure 7: Data collection in CREW. All human and agent data, including game states, feedback, mouse and keyboard, audio, gaze and pupil data, EEG and ECG data are all streamed and synchronized through Lab Streaming Layer.
@@ 3.6 Designing Algorithms
Algorithms research is crucial for Human-AI teaming. We designed the algorithm component of CREW to be extensible and accessible to the ML community. Algorithms are implemented in Python with PyTorch Paszke et al. (2019) and TorchRL Bou et al. (2023) which is part of the PyTorch ecosystem, making it easy for researchers to design and deploy new algorithms. We provide implementations of state-of-the-art human-guided machine learning algorithms and strong reinforcement learning baselines. The API for communication between a Unity instance and Python algorithm is implemented with MLAgents Juliani et al. (2018). All environments follow a uniform communication protocol for the observation and action data across tasks. Our modular design allows smooth switching between tasks and algorithms. Real human experiments are time-consuming and costly. To ease algorithm debugging, we implemented simulated human feedback providers as surrogates for real humans for complex tasks. These simulated feedback uses heuristics based on prior task knowledge and is not available in novel tasks in the real world; hence, they should be used solely as debugging tools, not as replacements for real human evaluations.